# presep= ]e1,e2,e3

### v1

@@ Appendix C Anti-clustering conditions for simulations
When checking that the time series in the simulation study satisfy the required conditions, Lemma 3 can used to verify anti-clustering conditions 2 and 6. In this appendix we prove that (11) holds for the proposed conditional Fréchet and Pareto models, and also provide a result that can be used to check the anti-clustering conditions for the proposed conditional max-stable process.
@@ C.1 Conditional Fréchet and Pareto
Consider covariate process $\left\{X_{j}\right\}$, belonging to some stationary (possibly long-memory) time series model and $\left\{U_{j}\right\}$ stationary (possibly long-memory) with marginals in $[0,1]$. The constructions
@@
[…]
@@
Let $s,t>0$ be given. In both (17) and (18), the Cauchy-Schwarz inequality yields that
$\frac{\mathbb{E}\left[1_{\left\{Y_{0}>su_{n}\right\}}1_{\left\{Y_{j}>tu_{n}\right\}}\mid X_{0}=x+h_{n}w_{1},X_{j}=x+h_{n}w_{2}\right]}{\overline{F}^{x}(u_{n})}$
$\leq$ $\frac{\mathbb{E}\left[1_{\left\{Y_{0}>su_{n}\right\}}\mid X_{0}=x+h_{n}w_{1},X_{j}=x+h_{n}w_{2}\right]^{1/2}\mathbb{E}\left[1_{\left\{Y_{j}>tu_{n}\right\}}\mid X_{0}=x+h_{n}w_{1},X_{j}=x+h_{n}w_{2}\right]^{1/2}}{\overline{F}^{x}(u_{n})}$
$=\left(\frac{\mathbb{E}\left[1_{\left\{Y_{0}>su_{n}\right\}}\mid X_{0}=x+h_{n}w_{1}\right]}{\overline{F}^{x}(u_{n})}\right)^{1/2}\left(\frac{\mathbb{E}\left[1_{\left\{Y_{j}>tu_{n}\right\}}\mid X_{j}=x+h_{n}w_{2}\right]}{\overline{F}^{x}(u_{n})}\right)^{1/2}=\mathcal{O}(1),$
uniformly in $w_{1},w_{2}\in[-1,1]$, where we have used Potter’s bound and the Lipschitz conditions on $L^{x}$ and $\alpha_{x}$ as was done in Lemma 4. This shows condition 6.
Showing (11) is done analogously
$\frac{\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}A}\right)\mid X_{0}=x+h_{n}w_{1},X_{j}=x+h_{n}w_{2}\right]}{\overline{F}^{x}(u_{n})}$
$\leq\left(\frac{\mathbb{E}\left[\log_{+}^{2}\left(\frac{Y_{0}}{u_{n}A}\right)\mid X_{0}=x+h_{n}w_{1}\right]}{\overline{F}^{x}(u_{n})}\right)^{1/2}\left(\frac{\mathbb{E}\left[\log_{+}^{2}\left(\frac{Y_{0}}{u_{n}A}\right)\mid X_{0}=x+h_{n}w_{2}\right]}{\overline{F}^{x}(u_{n})}\right)^{1/2}$
$\quad\to\int_{A}^{\infty}\log^{2}(t)\alpha_{x}t^{-\alpha_{x}-1}\mathrm{d}t,$
uniformly in $w_{1},w_{2}\in[-1,1]$ as shown in the proof of Theorem 5. We conclude that also (11) holds. ∎
@@ C.2 Conditional Max-Stable process
Figure 7: Diagram of the conditional Max-stable (CSGMS) construction. Since the Gaussian processes $\textbf{W}^{(i)},i\geq 1$ are all started in zero, the distribution of the processes at time $1$, namely $(W_{1}^{(i)})_{i=1}^{\infty}$, doesn’t depend on the starting values $(W_{0}^{(i)})_{i=1}^{\infty}$ as they are deterministic. Hence no arrow between $(W_{0}^{(i)})_{i=1}^{\infty}$ and $(W_{1}^{(i)})_{i=1}^{\infty}$.
@@
$\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>tu_{n}\right\}}\right]$
$=\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\mathbb{E}\left[1_{\left\{Y_{0}>su_{n}\right\}}\mid X_{0},X_{j}\right]\mathbb{E}\left[1_{\left\{Y_{j}>tu_{n}\right\}}\mid X_{0},X_{j}\right]\right]$
$=\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\mathbb{E}\left[1_{\left\{Y_{0}>su_{n}\right\}}\mid X_{0}\right]\mathbb{E}\left[1_{\left\{Y_{j}>tu_{n}\right\}}\mid X_{j}\right]\right]$
$=(h_{n}\overline{F}^{x}(u_{n}))^{2}\int_{[-1,1]^{2}}\biggl{\{}\frac{\mathbb{E}\left[1_{\left\{Y_{0}>su_{n}\right\}}\mid X_{0}=x+w_{1}h_{n}\right]}{\overline{F}^{x}(u_{n})}\frac{\mathbb{E}\left[1_{\left\{Y_{0}>tu_{n}\right\}}\mid X_{0}=x+w_{2}h_{n}\right]}{\overline{F}^{x}(u_{n})}$
×$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times g(x+w_{1}h_{n},x+w_{2}h_{n})K\left(w_{1}\right)K\left(w_{2}\right)\biggr{\}}\mathrm{d}(w_{1},w_{2})$
$=\mathcal{O}((h_{n}\overline{F}^{x}(u_{n}))^{2}),$
where we have used Potter’s bound and the Lipschitz conditions on $L^{x}$ and $\alpha_{x}$ as was done in Lemma 4. Consequently
$\lim_{m\to\infty}\limsup_{n\to\infty}\sum_{j=m}^{r_{n}}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>tu_{n}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n})}$
$\quad\leq\limsup_{n\to\infty}\sum_{j=1}^{r_{n}}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>tu_{n}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n})}$
$\quad\leq\limsup_{n\to\infty}\mathcal{O}\left(r_{n}h_{n}\overline{F}^{x}(u_{n})\right)=0,$
@@
$\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}A}\right)\right]$
$=(h_{n}\overline{F}^{x}(u_{n}))^{2}\int_{[-1,1]^{2}}\biggl{\{}\frac{\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)\mid X_{0}=x+h_{n}w_{1}\right]}{\overline{F}^{x}(u_{n})}\frac{\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)\mid X_{0}=x+h_{n}w_{2}\right]}{\overline{F}^{x}(u_{n})}$
×$\qquad\qquad\qquad\qquad\qquad\qquad\times g(x+w_{1}h_{n},x+w_{2}h_{n})K\left(w_{1}\right)K\left(w_{2}\right)\biggr{\}}\mathrm{d}(w_{1},w_{2})$
$=\mathcal{O}((h_{n}\overline{F}^{x}(u_{n}))^{2}),$
@@ Appendix D Kernel density estimation for mixing sequences
We begin with an auxiliary result which is in all subsequent proofs.
@@
Consider a stationary time series $\left\{\left(X_{j}\right),j\in\mathbb{Z}\right\}$ where each $X_{j}$ takes values in $[-1,1]$ and admits twice continuously differentiable density $g$. Let the kernel $K$ be a bounded, symmetric and defined on $[-1,1]$, and let the bandwidth sequence $\left\{h_{n}\right\}$ satisfy
$\lim_{n\to\infty}h_{n}=0,\quad\lim_{n\to\infty}nh_{n}=\infty.$
@@
The Taylor expansion $g(x+h_{n}u)=g(x)+h_{n}ug^{(1)}(x)+\frac{h_{n}^{2}u^{2}}{2}g^{(2)}(x)+\mathcal{O}\left(\lvert h_{n}u\rvert^{3}\right)$ yields
@@
Assume that $h_{n}\to 0$, $nh_{n}\to\infty$, that the time series $\left\{\left(X_{j}\right),j\in\mathbb{Z}\right\}$ is stationary with
@@
Similarly $\mathbb{E}\left[K^{2}\left(\frac{x-X_{0}}{h_{n}}\right)\right]=\mathcal{O}(h_{n}^{2})$, and consequently, for fixed $m\in\mathbb{N}$ we have
@@
Assume that $h_{n}\to 0$, $nh_{n}\to\infty$, that the time series $\left\{\left(X_{j}\right),j\in\mathbb{Z}\right\}$ is stationary, $\alpha$-mixing with $\alpha$-mixing coefficient bounded by $\mathcal{O}(j^{-\eta}),\eta>2$ for every $j\in\mathbb{Z}$, and each $X_{j}$ admits twice continuously differentiable density $g$, and the joint density $g_{X_{i},X_{j}}$ of $(X_{i},X_{j})$ exists, and is bounded for all $i,j\in\mathbb{Z}$. Then for any $x\in\mathbb{R}$, it holds that $g_{n}(x)\overset{\mathrm{Pr}}{\to}g(x)$.
@@
Assume that $h_{n}\to 0$, $nh_{n}\to\infty$, that the time series $\left\{\left(X_{j}\right),j\in\mathbb{Z}\right\}$ is stationary, $m$-dependent and each $X_{j}$ admits twice continuously differentiable density $g$. Then for any $x\in\mathbb{R}$, it holds that $g_{n}(x)\overset{\mathrm{Pr}}{\to}g(x)$.
@@
$\operatorname{var}\left(g_{n}^{(i)}(x)\right)$ $=\frac{m^{2}}{\left(nh_{n}\right)^{2}}\left\lfloor\frac{n}{m}\right\rfloor\operatorname{var}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)\right)=\frac{\frac{m}{n}\left\lfloor\frac{n}{m}\right\rfloor m}{nh_{n}}\left(\frac{\mathbb{E}\left[K^{2}\left(\frac{x-X_{0}}{h_{n}}\right)\right]}{h_{n}}-\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)\right]}{h_{n}}\right)\to 0,$
@@
Assume that that $h_{n}\to 0$, $u_{n}\to\infty$, $nh_{n}\to\infty$, that the time series $\left\{\left(X_{j}\right),j\in\mathbb{Z}\right\}$ is stationary, each $X_{j}$ admits twice continuously differentiable density $g$,
@@
$\lim_{m\to\infty}\limsup_{n\to\infty}\overline{F}^{x}(u_{n})\sum_{j=m}^{n}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\right]}{h_{n}}=0.$
@@
$\limsup_{n\to\infty}\frac{2\overline{F}^{x}(u_{n})}{h_{n}}\sum_{i=1}^{m}\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right),K\left(\frac{x-X_{i}}{h_{n}}\right)\right)\leq\text{cst}\limsup_{n\to\infty}m\overline{F}^{x}(u_{n})h_{n}=0,$
@@
$\limsup_{n\to\infty}\frac{2\overline{F}^{x}(u_{n})}{h_{n}}\sum_{i=m}^{n}\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right),K\left(\frac{x-X_{i}}{h_{n}}\right)\right)\leq\varepsilon,$
@@
Assume that (19) holds, $h_{n}\to 0$, $u_{n}\to\infty$, $nh_{n}\to\infty$, that the time series $\left\{\left(X_{j}\right),j\in\mathbb{Z}\right\}$ is stationary, $\alpha$-mixing with $\alpha$-mixing coefficient bounded by $\mathcal{O}(j^{-\eta}),\eta>2$ for every $j\in\mathbb{Z}$, and each $X_{j}$ admits twice continuously differentiable density $g$.
@@
$\text{var}\left(A_{n}\right)\leq\frac{\overline{F}^{x}(u_{n})}{h_{n}}\text{var}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)\right)+\frac{2\overline{F}^{x}(u_{n})}{h_{n}}\sum_{i=1}^{n}\left|\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right),K\left(\frac{x-X_{i}}{h_{n}}\right)\right)\right|.$
The first term has leading term of order $\overline{F}^{x}(u_{n})$, which vanishes asymptotically. For the second term, we again consider the integer sequence $\left\{d_{n},n\in\mathbb{N}\right\}$ given by $d_{n}=\lceil h_{n}^{-2/\eta}\rceil$. Using Davydov’s and Billingsley’s inequality with the $\alpha$-mixing assumption yields
$\frac{2\overline{F}^{x}(u_{n})}{h_{n}}\sum_{i=1}^{n}\left|\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right),K\left(\frac{x-X_{i}}{h_{n}}\right)\right)\right|\leq\text{cst}\frac{\overline{F}^{x}(u_{n})}{h_{n}}\left(\sum_{i=1}^{d_{n}-1}h_{n}^{2}+\sum_{i=d_{n}}^{\infty}\alpha(i)\right)$
$\leq\text{cst}\left(\overline{F}^{x}(u_{n})h_{n}d_{n}+\frac{\overline{F}^{x}(u_{n})}{h_{n}}d_{n}^{-\eta+1}\right)\sim\text{cst}\overline{F}^{x}(u_{n})h_{n}^{1-\frac{2}{\eta}}\to 0.$

### v2

@@ Appendix C Anti-clustering conditions for simulations
When checking that the time series in the simulation study satisfy the required conditions, Lemma 3 can used to verify anti-clustering conditions 2 and 6. In this appendix we prove that equation (8) holds for the proposed conditional Fréchet and Pareto models, and also provide a result that can be used to check the anti-clustering conditions for the proposed conditional max-stable process.
@@ C.1 Conditional Fréchet and Pareto
Consider covariate process $\{X_{j}\}$, belonging to some stationary (possibly long-memory) time series model and $\{U_{j}\}$ stationary (possibly long-memory) with marginals in $[0,1]$. The constructions
@@
gives rise to a time series models with marginal $\text{Fréchet}(1/\gamma(X_{j}),1,0)$ and $\text{Pareto}(1,1/\gamma(X_{j}))$ distributions respectively for every $j$.
@@
Let $s,t>0$ be given. In both equation (14) and (15), the Cauchy-Schwarz inequality yields that
$\mathbb{E}\left[1_{\left\{Y_{0}>su_{n}^{x}\right\}}1_{\left\{Y_{j}>tu_{n}^{x}\right\}}\mid X_{0}=x+h_{n}w_{1},X_{j}=x+h_{n}w_{2}\right]/\overline{F}^{x}(u_{n}^{x})$
$\leq$ $\frac{\mathbb{E}\left[1_{\left\{Y_{0}>su_{n}^{x}\right\}}\mid X_{0}=x+h_{n}w_{1},X_{j}=x+h_{n}w_{2}\right]^{1/2}\mathbb{E}\left[1_{\left\{Y_{j}>tu_{n}^{x}\right\}}\mid X_{0}=x+h_{n}w_{1},X_{j}=x+h_{n}w_{2}\right]^{1/2}}{\overline{F}^{x}(u_{n}^{x})}$
$=\left(\frac{\mathbb{E}\left[1_{\left\{Y_{0}>su_{n}^{x}\right\}}\mid X_{0}=x+h_{n}w_{1}\right]}{\overline{F}^{x}(u_{n}^{x})}\right)^{1/2}\left(\frac{\mathbb{E}\left[1_{\left\{Y_{j}>tu_{n}^{x}\right\}}\mid X_{j}=x+h_{n}w_{2}\right]}{\overline{F}^{x}(u_{n}^{x})}\right)^{1/2}=\mathcal{O}(1),$
uniformly in $w_{1},w_{2}\in[-1,1]$, where we have used Potter’s bound and the Lipschitz conditions on $L^{x}$ and $1/\gamma(x)$ as was done in Lemma 4. This shows condition 6.
Showing equation (8) is done analogously
$\frac{\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}^{x}A}\right)\mid X_{0}=x+h_{n}w_{1},X_{j}=x+h_{n}w_{2}\right]}{\overline{F}^{x}(u_{n}^{x})}$
$\leq\left(\frac{\mathbb{E}\left[\log_{+}^{2}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)\mid X_{0}=x+h_{n}w_{1}\right]}{\overline{F}^{x}(u_{n}^{x})}\right)^{1/2}\left(\frac{\mathbb{E}\left[\log_{+}^{2}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)\mid X_{0}=x+h_{n}w_{2}\right]}{\overline{F}^{x}(u_{n}^{x})}\right)^{1/2}$
$\quad\to\int_{A}^{\infty}\log^{2}(t)(1/\gamma(x))t^{-1/\gamma(x)-1}\mathrm{d}t,$
uniformly in $w_{1},w_{2}\in[-1,1]$ as shown in the proof of Theorem 5. We conclude that also equation (8) holds. ∎
@@ C.2 Conditional Max-Stable process
$X_{0}$$(P_{i}^{X_{0}})_{i=1}^{\infty}$$X_{1}$$(P_{i}^{X_{1}})_{i=1}^{\infty}$$X_{2}$$(P_{i}^{X_{2}})_{i=1}^{\infty}$$Y_{0}$$Y_{1}$$Y_{2}$$(W_{0}^{(i)})_{i=1}^{\infty}$$(W_{1}^{(i)})_{i=1}^{\infty}$$(W_{2}^{(i)})_{i=1}^{\infty}$
Figure 7: Diagram of the conditional Max-stable (CSGMS) construction. Since the Gaussian processes $\textbf{W}^{(i)},i\geq 1$ are all started in zero, the distribution of the processes at time 11, namely $(W_{1}^{(i)})_{i=1}^{\infty}$, doesn’t depend on the starting values $(W_{0}^{(i)})_{i=1}^{\infty}$ as they are deterministic. Hence no arrow between $(W_{0}^{(i)})_{i=1}^{\infty}$ and $(W_{1}^{(i)})_{i=1}^{\infty}$.
@@
$\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>tu_{n}^{x}\right\}}\right]$
$=\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\mathbb{E}\left[1_{\left\{Y_{0}>su_{n}^{x}\right\}}\mid X_{0},X_{j}\right]\mathbb{E}\left[1_{\left\{Y_{j}>tu_{n}^{x}\right\}}\mid X_{0},X_{j}\right]\right]$
$=\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\mathbb{E}\left[1_{\left\{Y_{0}>su_{n}^{x}\right\}}\mid X_{0}\right]\mathbb{E}\left[1_{\left\{Y_{j}>tu_{n}^{x}\right\}}\mid X_{j}\right]\right]$
$=(h_{n}\overline{F}^{x}(u_{n}^{x}))^{2}\int_{[-1,1]^{2}}\biggl\{\frac{\mathbb{E}\left[1_{\left\{Y_{0}>su_{n}^{x}\right\}}\mid X_{0}=x+w_{1}h_{n}\right]}{\overline{F}^{x}(u_{n}^{x})}\frac{\mathbb{E}\left[1_{\left\{Y_{0}>tu_{n}^{x}\right\}}\mid X_{0}=x+w_{2}h_{n}\right]}{\overline{F}^{x}(u_{n}^{x})}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times g(x+w_{1}h_{n},x+w_{2}h_{n})K\left(w_{1}\right)K\left(w_{2}\right)\biggr\}\mathrm{d}(w_{1},w_{2})$
$=\mathcal{O}((h_{n}\overline{F}^{x}(u_{n}^{x}))^{2}),$
where we have used Potter’s bound and the Lipschitz conditions on $L^{x}$ and $1/\gamma(x)$ as was done in Lemma 4. Consequently
$\lim_{m\to\infty}\limsup_{n\to\infty}\sum_{j=m}^{r_{n}}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>tu_{n}^{x}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n}^{x})}$
$\quad\leq\limsup_{n\to\infty}\sum_{j=1}^{r_{n}}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>tu_{n}^{x}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n}^{x})}\leq\limsup_{n\to\infty}\mathcal{O}\left(r_{n}h_{n}\overline{F}^{x}(u_{n}^{x})\right)=0,$
@@
$\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}^{x}A}\right)\right]$
$=(h_{n}\overline{F}^{x}(u_{n}^{x}))^{2}\int_{[-1,1]^{2}}\biggl\{\frac{\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)\mid X_{0}=x+h_{n}w_{1}\right]}{\overline{F}^{x}(u_{n}^{x})}\frac{\mathbb{E}\left[\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)\mid X_{0}=x+h_{n}w_{2}\right]}{\overline{F}^{x}(u_{n}^{x})}$
$\qquad\qquad\qquad\qquad\qquad\qquad\times g(x+w_{1}h_{n},x+w_{2}h_{n})K\left(w_{1}\right)K\left(w_{2}\right)\biggr\}\mathrm{d}(w_{1},w_{2})$
$=\mathcal{O}((h_{n}\overline{F}^{x}(u_{n}^{x}))^{2}),$
@@ Appendix D Kernel density estimation for mixing sequences
We begin with an auxiliary result which is used in all subsequent proofs.
@@
Consider a stationary time series $\{(X_{j}),j\in\mathbb{Z}\}$ where each $X_{j}$ takes values in [1,1][-1,1] and admits twice continuously differentiable density $g$. Let the kernel $K$ be a bounded, symmetric and defined on [1,1][-1,1], and let the bandwidth sequence $\left\{h_{n}\right\}$ satisfy
$\lim_{n\to\infty}h_{n}=0.$
@@
The Taylor expansion $g(x+h_{n}u)=g(x)+h_{n}ug^{(1)}(x)+h_{n}^{2}u^{2}g^{(2)}(x)/2+\mathcal{O}\left(\lvert h_{n}u\rvert^{3}\right)$ yields
@@
Assume that $h_{n}\to 0$, $nh_{n}\to\infty$, that the time series $\{(X_{j}),j\in\mathbb{Z}\}$ is stationary with
@@
Similarly $\mathbb{E}\left[K^{2}((x-X_{0})/h_{n})\right]=\mathcal{O}(h_{n}^{2})$, and consequently, for fixed $m\in\mathbb{N}$ we have
@@
Assume that $h_{n}\to 0$, $nh_{n}\to\infty$, that the time series $\{(X_{j}),j\in\mathbb{Z}\}$ is stationary, $\alpha$-mixing with $\alpha$-mixing coefficient bounded by $\mathcal{O}(j^{-\eta}),\eta>2$ for every $j\in\mathbb{Z}$, and each $X_{j}$ admits twice continuously differentiable density $g$, and the joint density $g_{X_{i},X_{j}}$ of $(X_{i},X_{j})$ exists, and is bounded for all $i,j\in\mathbb{Z}$. Then for any $x\in\mathbb{R}$, it holds that $g_{n}(x)\overset{\mathrm{Pr}}{\to}g(x)$.
@@
Assume that $h_{n}\to 0$, $nh_{n}\to\infty$, that the time series $\{(X_{j}),j\in\mathbb{Z}\}$ is stationary, $m$-dependent and each $X_{j}$ admits twice continuously differentiable density $g$. Then for any $x\in\mathbb{R}$, it holds that $g_{n}(x)\overset{\mathrm{Pr}}{\to}g(x)$.
@@
$\operatorname{var}\left(\frac{1}{m}\sum_{i=1}^{m}g_{n}^{(i)}(x)\right)$ $=\frac{1}{m}\operatorname{var}\left(g_{n}^{(i)}(x)\right)\frac{m}{\left(nh_{n}\right)^{2}}\left\lfloor\frac{n}{m}\right\rfloor\operatorname{var}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)\right)$
$=\frac{\frac{m}{n}\left\lfloor\frac{n}{m}\right\rfloor}{nh_{n}}\left(\frac{\mathbb{E}\left[K^{2}\left(\frac{x-X_{0}}{h_{n}}\right)\right]}{h_{n}}-\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)\right]}{h_{n}}\right)\to 0,$
@@
Assume that that $h_{n}\to 0$, $u_{n}^{x}\to\infty$, $nh_{n}\to\infty$, that the time series $\{(X_{j}),j\in\mathbb{Z}\}$ is stationary, each $X_{j}$ admits twice continuously differentiable density $g$,
@@
$\lim_{m\to\infty}\limsup_{n\to\infty}\overline{F}^{x}(u_{n}^{x})\sum_{j=m}^{n}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\right]}{h_{n}}=0.$
@@
$\limsup_{n\to\infty}\frac{2\overline{F}^{x}(u_{n}^{x})}{h_{n}}\sum_{i=1}^{m}\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right),K\left(\frac{x-X_{i}}{h_{n}}\right)\right)\leq\text{cst}\limsup_{n\to\infty}m\overline{F}^{x}(u_{n}^{x})h_{n}=0,$
@@
$\limsup_{n\to\infty}\frac{2\overline{F}^{x}(u_{n}^{x})}{h_{n}}\sum_{i=m}^{n}\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right),K\left(\frac{x-X_{i}}{h_{n}}\right)\right)\leq\varepsilon,$
@@
Assume that (16) holds, that $h_{n}\to 0$, $u_{n}^{x}\to\infty$, $nh_{n}\to\infty$, that the time series $\{(X_{j}),j\in\mathbb{Z}\}$ is stationary, $\alpha$-mixing with $\alpha$-mixing coefficient bounded by $\mathcal{O}(j^{-\eta}),\eta>2$ for every $j\in\mathbb{Z}$, and each $X_{j}$ admits twice continuously differentiable density $g$.
@@
$\text{var}\left(A_{n}\right)\leq\frac{\overline{F}^{x}(u_{n}^{x})}{h_{n}}\text{var}\left(K\left(\frac{x-X_{0}}{h_{n}}\right)\right)+\frac{2\overline{F}^{x}(u_{n}^{x})}{h_{n}}\sum_{i=1}^{n}\left|\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right),K\left(\frac{x-X_{i}}{h_{n}}\right)\right)\right|.$
The first term has leading term of order $\overline{F}^{x}(u_{n}^{x})$, which vanishes asymptotically. For the second term, we again consider the integer sequence $\left\{d_{n},n\in\mathbb{N}\right\}$ given by $d_{n}=\lceil h_{n}^{-2/\eta}\rceil$. Using Davydov’s and Billingsley’s inequality with the $\alpha$-mixing assumption yields
$\frac{2\overline{F}^{x}(u_{n}^{x})}{h_{n}}\sum_{i=1}^{n}\left|\text{cov}\left(K\left(\frac{x-X_{0}}{h_{n}}\right),K\left(\frac{x-X_{i}}{h_{n}}\right)\right)\right|\leq\text{cst}\frac{\overline{F}^{x}(u_{n}^{x})}{h_{n}}\left(\sum_{i=1}^{d_{n}-1}h_{n}^{2}+\sum_{i=d_{n}}^{\infty}\alpha(i)\right)$
$\leq\text{cst}\left(\overline{F}^{x}(u_{n}^{x})h_{n}d_{n}+\frac{\overline{F}^{x}(u_{n}^{x})}{h_{n}}d_{n}^{-\eta+1}\right)\sim\text{cst}\overline{F}^{x}(u_{n}^{x})h_{n}^{1-\frac{2}{\eta}}\to 0.$